OpenStudy (anonymous):

Find the slope of the line that passes through (3, -4) and (-1, 2).

OpenStudy (anonymous):

id like steps and explanations please

OpenStudy (anonymous):

Subtract the y values from each other and the x values from each other. Then divide y/x Ex. 2-(-4)=y (-1-3)=x y/x=slope

OpenStudy (anonymous):

-6/4=-3/2=slope

OpenStudy (anonymous):

hey the ans is 3/2

OpenStudy (anonymous):

2-(-2)= 4 and -1-3 equals -4 4/-4 that makes no sense

OpenStudy (anonymous):

sorry its -3/2

OpenStudy (anonymous):

i will explain

OpenStudy (anonymous):

i want to know HOW to get it so every time i have q i dont have to come back here

OpenStudy (anonymous):

let the (3,-4) be (x1,y1)

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

What? (2-(-4))=6 and (-1-3)=-4

OpenStudy (anonymous):

let (-1,2) be (x2,y2)

OpenStudy (anonymous):

You made the arithmetic error by using 2-(-2)= 4 . It's not -2 but -4.

OpenStudy (anonymous):

now u get x1=3, y1=-4, x2=-1, y2=2

OpenStudy (anonymous):

they just substitute the values in the formula y2-y1/x2-x1

OpenStudy (anonymous):

u will get the slope as -3/2

OpenStudy (anonymous):

tell me if u have got it

OpenStudy (anonymous):

um sorry i got lost at this part now u get x1=3, y1=-4, x2=-1, y2=2

OpenStudy (anonymous):

To find the slope of the line that passes through (3, -4) and (-1, 2): Use the simplified formula: m = (y2- y1) / (x2 - x1) m = 2 - (-4) / (-1 - 3) m = 6 / -4 m = -1.5 the slope of the line is -1.5